# ALPS II collaboration

### v1

@@ V WISP search results
The results of the ALPS II first science campaign are summarized here and compared to likewise model-independent or nearly model-independent outcomes from other experiments. A full review of the quickly evolving WISPy experimental landscape is beyond the scope of this paper and we refer the reader to reference [AxionLimits] for an overview. Table 4 summarizes the 95% CL upper limits for $F_{N,\Delta}\rightarrow 1$ resp. $F_{\gamma^{\prime}}\rightarrow 1$ (equations (4c), (6b)).
@@
For pseudoscalar bosons (Fig. 3) the sensitivity achieved by ALPS II is about a factor of 20 better than at previous LSW experiments. ALPS II aims for surpassing the limits derived from solar axion searches (set by CAST) and will be sensitive to more recent QCD axion models involving e.g. magnetic monopoles [Sokolov:2021ydn].
Figure 3: Limits on pseudoscalar bosons: “Previous laboratory” (grey) summarizes previous results from ALPS [Ehret:2010mh] and OSQAR [OSQAR:2015qdv]; CAST (light-blue) is taken from [CAST:2024eil] and references therein. The golden axion band shows an “artist’s” view on the approximate range given by KSVZ- and DFSZ-inspired models [PhysRevLett.43.103, SHIFMAN1980493, DINE1981199, Zhitnitsky:1980tq], while the yellow range refers to a more recent model [Sokolov:2021ydn]. The green area shows the result of this analysis; the purple line indicates the ALPS II prospects.
@@
The mixing of vector bosons with photons does not depend on the magnetic field. Hence, the results of the two data sets with the different polarizations ($\gamma_{\perp}$ and $\gamma_{\parallel}$) can be combined. We use a conservative approach by combining the limits on the power ratios of the closed- and open-shutter measurements $P_{\gamma}\,/\,P_{\mathrm{open}}$ (without the systematic uncertainties) via $\mathrm{{Limit(\perp+\parallel)=\left(Limit(\perp)^{-2}+Limit(\parallel)^{-2}\right)^{-0.5}}}$ (see [ParticleDataGroup:2024cfk], section “Statistics”). The systematic uncertainty is then added to the combined result. Note that in the first science campaign of ALPS II, $L_{\mathrm{osci}}=L_{\mathrm{V}}$} before the wall and $L_{\mathrm{osci}}=L_{\mathrm{RC}}$} behind it (eq. (6) and Tab. 2). As visible in Fig. 5, the improvement relative to earlier results is more modest, because the unique ALPS II magnet strings are irrelevant for vector boson searches.
Fig. 6 shows ALPS II exclusion limits on massive tensor bosons relative to the standard-graviton coupling strength $G=\mathrm{6.709\cdot 10^{-39}\,GeV^{-2}}$. Here, we combined the limits from the two polarization states like described above, taking into account the different factors in equations (8a) and (8b). Table 4 also lists the limits for the two tensor searches with different light polarizations. Note that in models involving Lorentz symmetry violation, tensor spin-0 polarization states do not propagate: in this case, the limits from fifth force experiments in Fig. 6 do not apply (see [zxtk-bwnf] for a discussion).
Figure 4: Limits on scalar bosons: “Fifth force” (tan) is taken from [Adelberger_2003, KONOPLIV2011401] with the scaling $\mathrm{}{Q_{e}\sim 1/500}$ used in [Arvanitaki_2016], “Previous laboratory” (grey) summarizes results from ALPS [Ehret:2010mh] and OSQAR [OSQAR:2015qdv]; “CAST” (light-blue) is taken from [CAST:2024eil] and references therein. The green area shows the result of this analysis; the purple line indicates the ALPS II prospects.
Figure 5: Limits on hidden photons: “Previous laboratory” (grey) summarizes results from CROWS [Betz_2013] and ALPS [Ehret:2010mh]; “CAST” data (light-blue) are taken from [Redondo_2008]. The green area shows the result of this analysis; the purple line indicates the ALPS II prospects.
Figure 6: Limits on tensor bosons coupling strength $G^{\prime}$ relative to the standard-graviton coupling $G$. ”Fifth force” (tan) is taken from [Cembranos:2017vgi]. The green area shows the result of this analysis; the purple line indicates the ALPS II prospects.

### v2

@@ V WISP search results
The results of the ALPS II first science campaign are summarized here and compared to likewise model-independent or nearly model-independent outcomes from other experiments. A full review of the quickly evolving WISPy experimental landscape is beyond the scope of this paper and we refer the reader to reference [45] for an overview. Table 4 summarizes the 95% CL upper limits for $F_{N,\Delta}\rightarrow 1$ resp. $F_{\gamma^{\prime}}\rightarrow 1$ (equations (4c), (6b)).
@@
For pseudoscalar bosons (Fig. 3) the sensitivity achieved by ALPS II is about a factor of 20 better than at previous LSW experiments. ALPS II aims for surpassing the limits derived from solar axion searches (set by CAST) and will be sensitive to more recent QCD axion models involving e.g. magnetic monopoles [61].
Figure 3: Limits on pseudoscalar bosons: “Previous laboratory” (grey) summarizes previous results from ALPS [25], OSQAR [11] and PVLAS [26]; CAST (light-blue) is taken from [5] and references therein. The golden axion band shows an “artist’s” view on the approximate range given by KSVZ- and DFSZ-inspired models [40, 57, 20, 65], while the yellow range refers to a more recent model [61]. The green area shows the result of this analysis; the purple line, the ALPS II prospects.
@@
The mixing of vector bosons with photons does not depend on the magnetic field. Hence, the results of the two data sets with the different polarizations ($\gamma_{\perp}$ and $\gamma_{\parallel}$) can be combined. We use a conservative approach by combining the limits on the power ratios of the closed- and open-shutter measurements $P_{\gamma}\,/\,P_{\mathrm{open}}$ (without the systematic uncertainties) via $\mathrm{{Limit(\perp+\parallel)=\left(Limit(\perp)^{-2}+Limit(\parallel)^{-2}\right)^{-0.5}}}$ (see [44], section “Statistics”). The systematic uncertainty is then added to the combined result. Note that in the first science campaign of ALPS II, $L_{\mathrm{osci}}=L_{\mathrm{V}}$} before the wall and $L_{\mathrm{osci}}=L_{\mathrm{RC}}$} behind it (eq. (6) and Tab. 2). As visible in Fig. 5, the improvement relative to earlier results is more modest, because the unique ALPS II magnet strings are irrelevant for vector boson searches.
Fig. 6 shows ALPS II exclusion limits on massive tensor bosons relative to the standard-graviton coupling strength $G=\mathrm{6.709\cdot 10^{-39}\,GeV^{-2}}$. Here, we combined the limits from the two polarization states like described above, taking into account the different factors in equations (8a) and (8b). Table 4 also lists the limits for the two tensor searches with different light polarizations. Note that in models involving Lorentz symmetry violation, tensor spin-0 polarization states do not propagate: in this case, the limits from fifth force experiments in Fig. 6 do not apply (see [13] for a discussion).
Figure 4: Limits on scalar bosons: “Fifth force” (tan) is taken from [3, 41] with the scaling $\mathrm{}{Q_{e}\sim 1/500}$ used in [9], “Previous laboratory” (grey) summarizes results from ALPS [25], OSQAR [11] and PVLAS [26]; “CAST” (light-blue) is taken from [5] and references therein. The green area shows the result of this analysis; the purple line, the ALPS II prospects.
Figure 5: Limits on hidden photons: “Previous laboratory” (grey) summarizes results from CROWS [12] and ALPS [25]; “CAST” data (light-blue) are taken from [52]. The green area shows the result of this analysis; the purple line, the ALPS II prospects.
Figure 6: Limits on tensor bosons coupling strength $G^{\prime}$ relative to the standard-graviton coupling $G$. ”Fifth force” (tan) is taken from [14]. The green area shows the result of this analysis; the purple line, the ALPS II prospects.